OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation 5x - 3y = -6

OpenStudy (rmrjr22):

y = mx + b ... does this look familiar?

OpenStudy (anonymous):

yes @rmrjr22

OpenStudy (rmrjr22):

good, now make ur equation look like it and you will get your answer

OpenStudy (texaschic101):

can you put the equation in y = mx + b form ?

OpenStudy (anonymous):

i think

OpenStudy (texaschic101):

lets try it......show me what you can do

OpenStudy (rmrjr22):

yeah, move this equation around!

OpenStudy (texaschic101):

5x - 3y = 6 you want to get y on one side

OpenStudy (anonymous):

-3y=5x-6

OpenStudy (rmrjr22):

keep going... get y all by itself

OpenStudy (texaschic101):

good start

OpenStudy (anonymous):

add 3 to both sides

OpenStudy (texaschic101):

no....divide both sides by -3

OpenStudy (rmrjr22):

nope... right now its multiplying with y... to get it to the other side you have to do the opposite

OpenStudy (anonymous):

oh thats what confused me

OpenStudy (rmrjr22):

watch ur signs btw

OpenStudy (anonymous):

so divide or multiply

OpenStudy (texaschic101):

divide

OpenStudy (anonymous):

ok

OpenStudy (rmrjr22):

when you have -5x +4 = 0 ... what do you do first?

OpenStudy (anonymous):

subtract 4

OpenStudy (texaschic101):

good

OpenStudy (rmrjr22):

then you have -5x = 4... now get x by its lonesome

OpenStudy (texaschic101):

you do the opposite

OpenStudy (anonymous):

wait i think i have the answer

OpenStudy (texaschic101):

-5x is multiplying, so do the opposite

OpenStudy (rmrjr22):

x = -4/5

OpenStudy (anonymous):

slope=5/3 y intercept=2

OpenStudy (texaschic101):

BINGO....see, you can do this

OpenStudy (anonymous):

ok thanks

OpenStudy (rmrjr22):

yup

OpenStudy (texaschic101):

5x - 3y = -6 -3y = -5x - 6 -3y/-3 = -5/3x - 6/-3 y = -5/3x + 2 slope = -5/3 and y intercept = 2

OpenStudy (texaschic101):

way to go @Courage. .....good job

OpenStudy (anonymous):

Thanks

OpenStudy (rmrjr22):

its 5/3... the negative cancels out